If a buffer solution is 0.120 M0.120 M in a weak acid (?a=7.5×10−5)Ka=7.5×10−5) and 0.440 M0.440 M in its conjugate base, what is the pH?
Answer #1

Answer

pH = 4.68

Explanation

pKa = - log Ka

pKa of weak acid = - log( 7.5 ×10-5) = 4.12

Henderson - Hasselbalch equation is

pH = pKa + log([A-]/[HA])

pH = 4.12 + log([0.440M/0.120M])

pH = 4.12 + 0.56

pH = 4.68